If X and Y are independent Poisson variates such that
P(X=1)=P(X=2) and P(Y=2)=P(Y=3).
Find the variance of X- 2Y.
1
Expert's answer
2020-03-09T13:51:25-0400

Let X follow Poisson("\lambda")

P(X=1) = "\frac{e^{-\lambda}\lambda^1}{1!}"


P(X=2) = "\frac{e^{-\lambda}\lambda^2}{2!}"


Using P(X =1)=P(X=2) we get "\fbox{$\lambda = 2$}"



Let Y follow Poisson("\lambda'" )

P(Y=2) = "\frac{e^{-\lambda'}\lambda'^2}{2!}"


P(Y=3) = "\\frac{e^{-\\lambda'}\\lambda'^3}{3!}\u200b"


Using P(Y =2)=P(Y=3) we get "\fbox{$\lambda' = 3$}"


Since, X and Y are independent random variables,

Var(X-2Y) = Var(X) + 4Var(Y) = "\lambda+4\times\lambda'"

= 2 + 4"\times"3 = 14


This is because Variance of a poisson random variable is its parameter itself